# The $L_{p}$-error rate for randomized quasi-Monte Carlo self-normalized importance sampling of unbounded integrands

### v1

@@ 2.1 Notation
In this work, all norms are taken to be Euclidean norms. To prevent confusion, we adopt the convention that bold symbols (e.g., $\bm{u}$, $\bm{v}$) denote vectors, while ordinary symbols denote scalars. For instance, $\bm{u}_{j}$ refers to the $j$-th point in a low-discrepancy point set, whereas $u_{j}$ denotes the $j$-th component of a $d$-dimensional vector $\bm{u}$. Denote $1\!:\!d={1,2,\ldots,d}$. Throughout, $\bm{v}$ will always denote a subset of 1:d1\!:\!d, $\bm{u}$ a point in $(0,1)^{d}$, and $\bm{x}$ a point in $\mathbb{R}^{d}$. The notation $\bm{a}_{\bm{v}}:\bm{b}_{\bar{\bm{v}}}$ denotes a vector in $\mathbb{R}^{d}$ whose $j$-th component is $a_{j}$ if $j\in\bm{v}$ and $b_{j}$ otherwise. And, $\partial^{\bm{v}}h(\bm{u})$ denotes the mixed partial derivative of $h(\bm{u})$ with respect to $u_{j}$ with $j\in\bm{v}$. For $\bm{a}=\left(a_{1},\ldots,a_{d}\right)\in\mathbb{N}_{0}^{d}$ where $\mathbb{N}_{0}^{d}$ is the set of all $d$-dimensional vectors of non-negative integers, define $|\bm{a}|=a_{1}+\cdots+a_{d},\ \bm{a}!=a_{1}!\ldots a_{d}!$, and
@@ 3 The $L_{p}$-error rate of RQMC
where $G_{i}$ denotes the CDF of a uni-variate distribution with density $g_{i}$, $i=1,2,\ldots,d$. General maps will be considered subsequently.
@@
For the function $\psi$ defined above, $\psi^{(k)}(0)=\psi^{(k)}(1)=0,1\leq k\leq d.$ Moreover, for $0<t<1$, there exists a constant $C>0$ such that $|\psi^{(k)}(t)|\leq C,1\leq k\leq d$.
@@
Denote $\zeta(s)=s^{d}(1-s)^{d}$ and $c=1/\int_{0}^{1}\zeta(s)ds=\frac{(2d+1)!}{(d!)^{2}}$, so that $\psi(t)=c\int_{0}^{t}\zeta(s)ds$. By the Leibnitz formula, $\psi^{(k)}(t)=c\zeta^{(k-1)}(t),1\leq k\leq d$, where
$\zeta^{(k-1)}(t)$ $=\sum_{i=0}^{k-1}\binom{k-1}{i}(t^{d})^{(i)}((1-t)^{d})^{(k-1-i)}.$
@@
$\psi^{(k)}(0)=c\zeta^{(k-1)}(0)=0,\quad\psi^{(k)}(1)=c\zeta^{(k-1)}(1)=0.$
Moreover, for $0\leq t\leq 1$, each term $(t^{d})^{(i)},((1-t)^{d})^{(k-1-i)}$ is bounded by a constant depending only on $d$ and $k$. Hence, for $1\leq k\leq d$, there exists a constant $C>0$ such that
$|\psi^{(k)}(t)|=|c\zeta^{(k-1)}(t)|\leq C.$
@@
A comparison of the two projection operators is given in Figure 1, where we illustrate the first component of the two-dimensional projection operator with $d=2,r=10,\delta=0.1$. It can be observed that, in our operator (10), the function value of the operator depends on both components, whereas the operator of [ouyang2024] depends only on the first component.
@@
    $\partial^{\bm{v}}P_{r}^{(i)}(\bm{x})$ exists for any $\bm{x}\in\mathbb{R}^{d}$.
@@
    $\left|\partial^{\bm{v}}P_{r}^{(i)}(\bm{x})\right|\lesssim\mathds{1}_{\{\|\bm{x}\|<r\}}$.
@@
Using the new projection operator (10), for transport map $\mathcal{T}$ satisfying (7) and $\{\bm{u}_{1},\ldots,\bm{u}_{N}\}$ satisfying (6), one of our main results is following.
@@
$|\partial^{\bm{v}}h(\bm{u})|\lesssim\exp\left(M(1-\delta/2)^{2}r^{2}\right)\mathds{1}_{\{\|\bm{x}\|<r\}}\prod_{j\in\bm{v}}\frac{1}{g_{j}(G_{j}^{-1}(u_{j}))},$
@@
In the last inequality, the spherical coordinate transformation is used. Hence, by the definition of the variation in the sense of Hardy and Krause [basu2016], we have
@@
Assume that $f$ satisfies Assumption 1 with $M_{f}\in\mathbb{R}$, and that $\tau$ satisfies Assumption 3 with $C_{\tau},M_{\tau}\in\mathbb{R}$, then for any $\varepsilon>0$, $f_{\tau}=f\circ\tau$ satisfies Assumption 1 with a growth rate $M_{f,\tau}$, where
$M_{f,\tau}:=M_{f}C_{\tau}^{2}(1+\varepsilon)+dM_{\tau}$ (14)
@@
Now, we consider the specific forms of the transformation $\tau$. For convenience, we first consider $\tau(\bm{x})=T(L\bm{x}+\bm{\mu})$, where $T(\bm{x})=\left(T_{1}(x_{1}),T_{2}(x_{2}),\ldots,T_{d}(x_{d})\right)$ is a component-wise nonlinear transformation with each $T_{i}$ being a nonlinear function, $L\in\mathbb{R}^{d\times d}$ is a lower-triangular matrix (scaled transformation), and $\bm{\mu}\in\mathbb{R}^{d}$ (shifted transformation).
@@
$|T_{i}(y)|\leq C_{T}|y|+C_{1},\quad|T_{i}^{(m)}(y)|\lesssim\exp(M_{T}|y|^{2}),\quad\forall y\in\mathbb{R},$ (16)
@@
$D^{\bm{a}}\tau_{i}(\bm{x})=T_{i}^{(|\bm{a}|)}(y_{i})\prod_{j=1}^{d}(L_{i,j})^{a_{j}},$
where $y_{i}=L_{i}\bm{x}+\mu_{i}$. Since $|T_{i}^{(m)}(y)|\lesssim\exp(M_{T}|y|^{2})$ for any $1\leq m\leq d$, it follows that for any $\varepsilon>0$,
@@
Consider the case of $T$ satisfying (16) and the standard Gaussian as the base distribution. Let $\bm{X}=\mathcal{T}(\bm{u})=T(\bm{y})$ where $\bm{y}=L\Phi^{-1}(\bm{u})+\bm{\mu}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})$ and $\bm{\Sigma}=L^{T}L$. For any $\varepsilon>0$, $\bm{X}$ satisfies Assumption 2 with $\eta=d-2$ and $\alpha=\frac{\lambda_{\min}(\bm{\Sigma}^{-1})}{2C_{T}^{2}}-\varepsilon$. Indeed, note that $\|\bm{X}\|\leq C_{T}\|\bm{y}\|+\sqrt{d}C_{1}$. Then, for any $t>C_{T}/\sqrt{\lambda_{\min}(\bm{\Sigma}^{-1})}+\sqrt{d}C$ and $\varepsilon>0$, it follows that
@@
It is straightforward to check that the following continuous nonlinear functions $T_{i}$ commonly used as the activation function in machine learning [sharma2017activation] satisfy (16).
@@
$D^{\bm{\lambda}}H_{i}(\tilde{\bm{x}})=(T_{i}^{K})^{(|\bm{\lambda}|)}(\mathbf{y})\prod_{j=1}^{d}(L_{i,j})^{\lambda_{j}},$
where $\mathbf{y}=L_{i}^{K}\tilde{\bm{x}}+{\mu}_{i}^{K}$, Hence,
@@ 4 The $L_{p}$-error rate of RQMC-based SNIS estimator
Assume that $1/q$, $\bar{\pi}$, and $f$ each satisfy Assumption 1 with respective growth rates $M_{q},~M_{\pi}$, and $M_{f}\in\mathbb{R}$. Then $\omega$ and $\omega f$ each satisfy Assumption 1 with respective growth rates $M_{\omega}=M_{q}+M_{\pi}$ and $M_{\omega f}=M_{q}+M_{\pi}+M_{f}$.
@@
Assume that the transport map $\mathcal{T}=\tau\circ G^{-1}$ satisfies Assumption 2 with $\alpha>0$ and $\eta\in\mathbb{R}$, and that transformation $\tau$ satisfies Assumption 3 with $C_{\tau},M_{\tau}\in\mathbb{R}$. Suppose that $1/q$, $\bar{\pi}$, and $f$ each satisfy Assumption 1 with respective growth rates $M_{q},M_{\pi}\in\mathbb{R}$, and $M_{f}>0$. If $M_{\omega f,\tau}<\alpha/p$, then for any $\epsilon>0$, we have
@@
Consider the setting of Corollary 3. Assume that $M_{\omega,\tau}>0$ is arbitrarily small and $M_{f}>0$. Suppose that there exists a constant $v>1$ such that $q\left(|f|^{v}\right)=\int_{\mathbb{R}^{d}}|f(\bm{x})|^{v}q(\bm{x})d\bm{x}<\infty$. For any $\epsilon>0$ and $1\leq p<\min\left\{v,\frac{\alpha}{M_{\omega f,\tau}}\right\},$ then the RQMC-SNIS estimator satisfies
$\left\{\mathbb{E}\left|\frac{q_{N}(\omega f)}{q_{N}(\omega)}-\frac{q(\omega f)}{q(\omega)}\right|^{p}\right\}^{1/p}=\mathcal{O}\left(N^{-\beta+\epsilon}\right),$
where $\beta=1-pM_{\omega f,\tau}/\alpha.$ Moreover, if $M_{\omega f,\tau}<\alpha/(2p)$, then $\beta>1/2$.
@@
$\mathbb{E}\left[\left|\frac{q_{N}(f\omega)}{q_{N}(\omega)}-\frac{q(\omega f)}{q(\omega)}\right|^{p}\right]=\mathbb{E}\left[\left|\frac{q_{N}(f\omega)}{q_{N}(\omega)}-\frac{q(\omega f)}{q(\omega)}\right|^{p}\mathds{1}_{\mathcal{E}}\right]+\mathbb{E}\left[\left|\frac{q_{N}(f\omega)}{q_{N}(\omega)}-\frac{q(\omega f)}{q(\omega)}\right|^{p}\mathds{1}_{\mathcal{E}^{c}}\right].$ (20)
@@
For any $1\leq p\leq\dfrac{sv}{s+v+1/(1-\epsilon)}$, we have
$p/v-(1-\epsilon)s(1-p/v)\leq-(1-\epsilon)p.$ (21)
Note that $\frac{sv}{s+v+1/(1-\epsilon)}$ is monotonically increasing for $s>0$ and $\lim\limits_{s\to\infty}\frac{sv}{s+v+1/(1-\epsilon)}=~v$. Since $M_{\omega,\tau}$ is arbitrarily small, $s$ can be chosen sufficiently large such that if $1\leq p<v$, then (21) holds. Moreover, note that $\frac{sv}{s+v+1/(1-\epsilon)}$ is also monotonically increasing for $v>0$ and $\lim\limits_{v\to\infty}\frac{sv}{s+v+1/(1-\epsilon)}=s$. Then, for any $s\geq 1$, we have
@@
Corollary 4.
An unbounded function $f$ that satisfies Assumption 2 with an arbitrarily small growth rate $M>0$ is referred to as a “QMC-friendly” function [he2024]. If function $\omega f$ is QMC-friendly, then $M_{\omega f}>0$ is arbitrarily small. When $M_{\tau}=0$, $M_{\omega f,\tau}=M_{\omega f}C_{\tau}^{2}(1+\varepsilon)+dM_{\tau}=M_{\omega f}C_{\tau}^{2}(1+\varepsilon)$ remains arbitrarily small. Since $M_{\omega f}=M_{\omega}+M_{f}$ with $M_{f}>0$, we have $M_{\omega}<M_{\omega f}$ and $M_{\omega,\tau}<M_{\omega f,\tau}$, and thus $M_{\omega,\tau}$ is arbitrarily small. Under the setting of Theorem 3, for any $\epsilon>0$ and $1\leq p<v$, we have $\left\{\mathbb{E}\left|\frac{q_{N}(\omega f)}{q_{N}(\omega)}-\frac{q(\omega f)}{q(\omega)}\right|^{p}\right\}^{1/p}=\mathcal{O}\left(N^{-1+\epsilon}\right).$
@@
Consider the setting of Corollary 4. Compared with the $L_{p}$-error rate $\mathcal{O}(N^{-1/2})$ in Theorem 2.3 of [deligiannidis2024], our result achieves a faster rate. Moreover, we require weaker moment condition on $f$ and $\omega$. In Theorem 2.3 of [deligiannidis2024], for constant $s\geq 2$, it needs $q(\omega^{s})<\infty$ and the condition on $p$ is $2\leq p\leq\frac{sv}{s+v+2}$. For instance, consider the case of $p=2$ (RMSE) and $s=3$, [deligiannidis2024] needs $v=10$ to achieve an error rate of $\mathcal{O}(N^{-1/2})$. In contrast, by Corollary 4, our result only requires $v=2+\delta$ to achieve an error rate of $\mathcal{O}(N^{-1+\epsilon})$ for any $\delta>0$ and $\epsilon>0$.

### v2

@@ 2.1 Notation
In this work, all norms are taken to be Euclidean norms. To prevent confusion, we adopt the convention that bold symbols (e.g., $\bm{u}$, $\bm{v}$) denote vectors, while ordinary symbols denote scalars. For instance, $\bm{u}_{j}$ refers to the $j$-th point in a low-discrepancy point set, whereas $u_{j}$ denotes the $j$-th component of a $d$-dimensional vector $\bm{u}$. Denote $1\!:\!d={1,2,\ldots,d}$. Throughout, $\bm{v}$ will always denote a subset of 1:d1\!:\!d, $\bm{u}$ a point in $(0,1)^{d}$, and $\bm{x}$ a point in $\mathbb{R}^{d}$. The notation $\bm{a}_{\bm{v}}:\bm{b}_{\bar{\bm{v}}}$ denotes a vector in $\mathbb{R}^{d}$ whose $j$-th component is $a_{j}$ if $j\in\bm{v}$ and $b_{j}$ otherwise. And, $\partial^{\bm{v}}h(\bm{u})$ denotes the mixed partial derivative of $h(\bm{u})$ with respect to $u_{j}$ with $j\in\bm{v}$. For $\bm{a}=\left(a_{1},\ldots,a_{d}\right)\in\mathbb{N}_{0}^{d}$ where $\mathbb{N}_{0}^{d}$ is the set of all $d$-dimensional vectors of non-negative integers, define $|\bm{a}|=a_{1}+\cdots+a_{d}$, and
@@ 3 The $L_{p}$-error rate of RQMC
where $G_{i}$ denotes the CDF of a uni-variate distribution with density $g_{i}$, $i=1,2,\ldots,d$. General transport maps will be considered subsequently.
@@
For the function $\psi$ defined in (10), $D^{k}\psi(0)=D^{k}\psi(1)=0,1\leq k\leq d.$ Moreover, for $0<t<1$, there exists a constant $C>0$ such that $|D^{k}\psi(t)|\leq C,1\leq k\leq d$.
@@
Denote $\zeta(s)=s^{d}(1-s)^{d}$ and $c=1/\int_{0}^{1}\zeta(s)ds=\frac{(2d+1)!}{(d!)^{2}}$, so that $\psi(t)=c\int_{0}^{t}\zeta(s)ds$. By the Leibnitz formula, $D^{k}\psi(t)=cD^{k-1}\zeta(t),1\leq k\leq d$, where
$D^{k-1}\zeta(t)$ $=\sum_{i=0}^{k-1}\binom{k-1}{i}D^{i}(t^{d})D^{k-1-i}((1-t)^{d}).$
@@
$D^{k}\psi(0)=cD^{k-1}\zeta(0)=0,\quad D^{k}\psi(1)=cD^{k-1}\zeta(1)=0.$
Moreover, for $0\leq t\leq 1$, each term $D^{i}(t^{d}),D^{k-1-i}((1-t)^{d})$ is bounded by a constant depending only on $d$ and $k$. Hence, for $1\leq k\leq d$, there exists a constant $C>0$ such that
$|D^{k}\psi(t)|=|cD^{k-1}\zeta(t)|\leq C.$
@@
A comparison of the two projection operators is given in Figure 1, where we illustrate the first component $P_{r}^{(1)}(\bm{x})$ of the two-dimensional projection operator with $d=2,r=10,\delta=0.1$. It can be observed that, in our operator (11), the function value of the operator depends on both components, whereas the operator of [ouyang2024] depends only on the first component.
@@
    $\partial^{\bm{v}}P_{r}^{(i)}(\bm{x})$ exists for any $\bm{x}\in\mathbb{R}^{d}$,
@@
    $\left|\partial^{\bm{v}}P_{r}^{(i)}(\bm{x})\right|\lesssim\mathds{1}_{\{\|\bm{x}\|<r\}}$,
@@
Using the new projection operator (11), for the transport map $\mathcal{T}$ satisfying (7) and $\{\bm{u}_{1},\ldots,\bm{u}_{N}\}$ satisfying (6), one of our main results is following.
@@
$|\partial^{\bm{v}}h(\bm{u})|\lesssim\exp\left(M(1-\delta/2)^{2}r^{2}\right)\mathds{1}_{\{\|\bm{x}\|<r\}}\prod_{j\in\bm{v}}\frac{1}{g_{j}(G_{j}^{-1}(u_{j}))}.$
@@
In the last inequality, the spherical coordinate transformation was used. Hence, by the definition of the variation in the sense of Hardy and Krause [basu2016], we have
@@
Assume that $f$ satisfies Assumption 1 with $M_{f}\in\mathbb{R}$, and that $\tau$ satisfies Assumption 3 with $C_{\tau},M_{\tau}\in\mathbb{R}$, then for any $\varepsilon>0$, $f_{\tau}=f\circ\tau$ satisfies Assumption 1 with a growth rate
$M_{f,\tau}=M_{f}C_{\tau}^{2}(1+\varepsilon)+dM_{\tau}.$ (15)
@@
Now, we consider a specific form of the transformation $\tau(\bm{x})=T(L\bm{x}+\bm{\mu})$, where $T(\bm{x})=\left(T_{1}(x_{1}),T_{2}(x_{2}),\ldots,T_{d}(x_{d})\right)$ is a component-wise nonlinear transformation with each $T_{i}$ being a nonlinear function, $L\in\mathbb{R}^{d\times d}$ is a lower-triangular matrix (scaled transformation), and $\bm{\mu}\in\mathbb{R}^{d}$ (shifted transformation).
@@
$|T_{i}(y)|\leq C_{T}|y|+C_{1},\quad|D^{m}T_{i}(y)|\lesssim\exp(M_{T}|y|^{2}),\quad\forall y\in\mathbb{R},$ (17)
@@
$D^{\bm{a}}\tau_{i}(\bm{x})=D^{|\bm{a}|}T_{i}(y_{i})\prod_{j=1}^{d}(L_{i,j})^{a_{j}},$
where $y_{i}=L_{i}\bm{x}+\mu_{i}$. Since $|D^{m}T_{i}(y)|\lesssim\exp(M_{T}|y|^{2})$ for any $1\leq m\leq d$, it follows that for any $\varepsilon>0$,
@@
Consider the case of $T$ satisfying (17) and the standard Gaussian as the base distribution. Let $\bm{X}=\mathcal{T}(\bm{u})=T(\bm{y})$ where $\bm{y}=L\Phi^{-1}(\bm{u})+\bm{\mu}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})$ and $\bm{\Sigma}=L^{T}L$. For any $\varepsilon>0$, $\bm{X}$ satisfies Assumption 2 with $\eta=d-2$ and $\alpha=\lambda_{\min}(\bm{\Sigma}^{-1})/(2C_{T}^{2})-\varepsilon$. Indeed, note that $\|\bm{X}\|\leq C_{T}\|\bm{y}\|+\sqrt{d}C_{1}$. Then, for any $t>C_{T}/\sqrt{\lambda_{\min}(\bm{\Sigma}^{-1})}+\sqrt{d}C$ and $\varepsilon>0$, it follows that
@@
It is straightforward to check that the continuous nonlinear functions $T_{i}$ in Table 1 commonly used as the activation function in machine learning [sharma2017activation] satisfy (17).
@@
$D^{\bm{\lambda}}H_{i}(\tilde{\bm{x}})=D^{|\bm{\lambda}|}T_{i}^{K}(y_{i})\prod_{j=1}^{d}(L_{i,j})^{\lambda_{j}},$
where $y_{i}=L_{i}^{K}\tilde{\bm{x}}+{\mu}_{i}^{K}$, Hence,
@@ 4 The $L_{p}$-error rate of RQMC-based SNIS estimator
Assume that $1/q$, $\bar{\pi}$, and $f$ each satisfy Assumption 1 with growth rates $M_{q},~M_{\pi}$, and $M_{f}\in\mathbb{R}$, respectively. Then $\omega$ and $\omega f$ each satisfy Assumption 1 with respective growth rates $M_{\omega}=M_{q}+M_{\pi}$ and $M_{\omega f}=M_{q}+M_{\pi}+M_{f}$.
@@
Assume that the transport map $\mathcal{T}=\tau\circ G^{-1}$ satisfies Assumption 2 with $\alpha>0$ and $\eta\in\mathbb{R}$, and that transformation $\tau$ satisfies Assumption 3 with $C_{\tau},M_{\tau}\in\mathbb{R}$. Suppose that $1/q$, $\bar{\pi}$, and $f$ satisfy Assumption 1 with respective growth rates $M_{q},M_{\pi}\in\mathbb{R}$, and $M_{f}>0$. If $M_{\omega f,\tau}<\alpha/p$, then for any $\epsilon>0$, we have
@@
Consider the setting of Corollary 3. Assume that $M_{\omega,\tau}>0$ is arbitrarily small and $M_{f}>0$. Suppose that there exists a constant $v>1$ such that $q\left(|f|^{v}\right)=\int_{\mathbb{R}^{d}}|f(\bm{x})|^{v}q(\bm{x})d\bm{x}<\infty$. If $1\leq p<\min\left\{v,\alpha/M_{\omega f,\tau}\right\}$, then for any $\epsilon>0$, the RQMC-SNIS estimator has a $L_{p}$-error rate
$\left\{\mathbb{E}\left|\pi_{N}(f)-\pi(f)\right|^{p}\right\}^{1/p}=\mathcal{O}\left(N^{-\beta+\epsilon}\right),$
where $\beta=1-pM_{\omega f,\tau}/\alpha>0.$ Moreover, if $M_{\omega f,\tau}<\alpha/(2p)$, then $\beta>1/2$.
@@
$\mathbb{E}\left[\left|\pi_{N}(f)-\pi(f)\right|^{p}\right]=\mathbb{E}\left[\left|\frac{q_{N}(f\omega)}{q_{N}(\omega)}-\frac{q(\omega f)}{q(\omega)}\right|^{p}\mathds{1}_{\mathcal{E}}\right]+\mathbb{E}\left[\left|\frac{q_{N}(f\omega)}{q_{N}(\omega)}-\frac{q(\omega f)}{q(\omega)}\right|^{p}\mathds{1}_{\mathcal{E}^{c}}\right].$ (21)
@@
To ensure
$p/v-(1-\epsilon)s(1-p/v)\leq-(1-\epsilon)p,$ (22)
it is required that $1\leq p\leq\dfrac{sv}{s+v+1/(1-\epsilon)}$. Note that $\frac{sv}{s+v+1/(1-\epsilon)}$ is monotonically increasing for $s>0$ and
$\lim\limits_{s\to\infty}\frac{sv}{s+v+1/(1-\epsilon)}=v.$
Since $M_{\omega,\tau}>0$ is arbitrarily small, we take a sufficiently large $s$ such that (22) holds for any $1\leq p<v$. It then follows (22) that for any $s\geq 1$,
@@
An unbounded function $f$ that satisfies Assumption 2 with an arbitrarily small growth rate $M>0$ is referred to as a “QMC-friendly” function [he2024]. If function $\omega f$ is QMC-friendly, then $M_{\omega f}>0$ is arbitrarily small. When $M_{\tau}=0$, $M_{\omega f,\tau}=M_{\omega f}C_{\tau}^{2}(1+\varepsilon)+dM_{\tau}=M_{\omega f}C_{\tau}^{2}(1+\varepsilon)$ remains arbitrarily small. Since $M_{\omega f}=M_{\omega}+M_{f}$ with $M_{f}>0$, we have $M_{\omega}<M_{\omega f}$ and $M_{\omega,\tau}<M_{\omega f,\tau}$, and thus $M_{\omega,\tau}$ is arbitrarily small. By Theorem 3, for any $\epsilon>0$ and $1\leq p<v$, we have $\left\{\mathbb{E}\left|\pi_{N}(f)-\pi(f)\right|^{p}\right\}^{1/p}=\mathcal{O}\left(N^{-1+\epsilon}\right).$ Compared with the $L_{p}$-error rate $\mathcal{O}(N^{-1/2})$ in Theorem 2.3 of [deligiannidis2024], our result achieves a faster rate. Moreover, we require weaker moment condition on $f$ and $\omega$. In Theorem 2.3 of [deligiannidis2024], for constant $s\geq 2$, it needs $q(\omega^{s})<\infty$ and the condition on $p$ is $2\leq p\leq\frac{sv}{s+v+2}$. For instance, consider the case of $p=2$ (RMSE) and $s=3$, [deligiannidis2024] needs $v=10$ to achieve an error rate of $\mathcal{O}(N^{-1/2})$. In contrast, if function $\omega f$ is QMC-friendly, our result only requires $v=2+\delta$ to achieve an error rate of $\mathcal{O}(N^{-1+\epsilon})$ for any $\delta>0$ and $\epsilon>0$.